8x+5y=11 and 3x-4y=10
Answer by Cromlix(4381)

Hi there,
8x + 5y = 11.....(1)
3x - 4y = 10.....(2)
Multiply (1) by 4 and
(2) by 5
32x + 20y = 44....(1)
15x - 20y = 50....(20
Add (1) + (2)
47x = 94
x = 2
Substitute x = 2 into
Equation (1)
8x + 5y = 11
8(2) + 5y = 11
16 + 5y = 11
5y = 11 - 16
5y = -5
y = -1
x = 2: y = -1
Hope this helps:-)